The following refer to the following data set:

57.6 55.8 41.2 77.9 77.9
53.5 77.9 45.5 30.6 62.5



What is the arithmetic mean of this data set?

mean =

What is the median of this data set?

median =

What is the mode of this data set?

mode =
Answer #1

a)

Mean = sum/n =  58.04

b)

Median = (55.8 + 57.6)/2 = 56.7

30.6
41.2
45.5
53.5
55.8
57.6
62.5
77.9
77.9
77.9

c)

Mode = 77.9